OpenStudy (elenathehomeschooler):

Can someone please explain how to solve this! simplify the expressions using imaginary numbers sqrt. -16

OpenStudy (amistre64):

-16 = -1*16 right?

OpenStudy (elenathehomeschooler):

Yes

OpenStudy (amistre64):

then split the radical and work the definitions ...

OpenStudy (amistre64):

\[\sqrt{-16}\] \[\sqrt{-1*16}\] \[\sqrt{-1}~\sqrt{16}\]

OpenStudy (elenathehomeschooler):

would that be the answer?

OpenStudy (elenathehomeschooler):

@amistre64 ?

OpenStudy (amistre64):

can sqrt(16) be simplified? can sqrt(-1) be simplified? or restated?

OpenStudy (amistre64):

sites getting laggy, at least on my end ...

OpenStudy (elenathehomeschooler):

the sqrt of 16 is 4 right?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

and by definintoin, i = sqrt(-1)

OpenStudy (elenathehomeschooler):

is the answer 

OpenStudy (amistre64):

no, sqrt(16) is not sqrt(4) ..... 4i is fine

OpenStudy (elenathehomeschooler):

oh okay thanks !

OpenStudy (amistre64):

youre welcome